# PositionIC: Unified Position and Identity Consistency for Image Customization

### v3

@@ Abstract
Recent subject-driven image customization has achieved significant advancements in fidelity, yet fine-grained instance-level spatial control remains elusive, hindering the broader real-world application. This limitation is mainly attributed to the absence of scalable datasets that bind identity with precise positional cues. To this end, we introduce PositionIC, a unified framework that enforces position and identity consistency for multi-subject customization. We construct a scalable synthesis pipeline that employs a bidirectional generation paradigm to eliminate subject drift and maintain semantic coherence. On top of these data, we design a lightweight positional modulation operation that decouples spatial embeddings among subjects, enabling independent, accurate placement while preserving visual fidelity. Extensive experiments demonstrate that our approach can achieve precise spatial control while maintaining high consistency in image customization task. PositionIC paves the way for controllable, high-fidelity image customization in open-world, multi-entity scenarios and will be released to foster further research.
Introduction
Diffusion-based models have recently revolutionized visual synthesis, especially in text-to-image (T2I) generation, where they produce photorealistic images that faithfully match short textual prompts (Podell et al. 2023; Esser et al. 2024; Rombach et al. 2022; Li et al. 2024; Labs 2024).
Within this broader area, subject-driven image customization seeks to generate new scenes that simultaneously (i) conform to user-provided textual descriptions and (ii) preserve the identity of one or more reference objects. Although recent methods (Wu et al. 2025; Mou et al. 2025; Tan et al. 2024; Wang et al. 2025) markedly improve visual fidelity, they still offer limited control over where or how each subject appears (i.e., position, scale, and relative layout). This limitation stems from twofold root causes: scarce data with explicit positional labels and existing global-level attention mechanisms that cannot deliver fine-grained, instance-level guidance.
@@ Introduction
Approaches that do permit object placement (Wang et al. 2024c; Li et al. 2023) often compromise either spatial accuracy or identity consistency, restricting their usefulness in practical scenarios such as e-commerce product display, storybook illustration, and interior design.
Consequently, it remains an open challenge to achieve both high-fidelity rendering and fine-grained, flexible control over subject attributes in a unified framework.
To this end, we address these issues with two tightly coupled innovations: i) Bidirectional Multi-dimensional Perception Data Synthesis (BMPDS), a scalable pipeline for generating high-quality, position-annotated multi-subject data; ii) PositionIC, a lightweight, layout-aware framework that enables fine-grained, position-controllable, identity-consistent image customization.
Specifically, to overcome the data bottleneck, we devise an automatic pipeline that expands single-subject collections into scalable, high-quality multi-subject datasets annotated with explicit position masks. A hierarchical training schedule establishes a bidirectional generation paradigm, progressively moving from single-to-multi subject synthesis and back, so that resolution constraints are relaxed while subject drift is suppressed. Because of the inherent hallucination of Multi-modal Large Language Models (MLLMs) (Han et al. 2024; Pi et al. 2024), we avoid direct visual comparisons: expert vision models first translate visual content into textual descriptions, after which MLLMs perform multi-dimensional consistency checks. This two-stage filtering markedly improves data reliability.
Built upon the BMPDS corpus, PositionIC injects an instance-level attention modulation into Diffusion Transformer (DiT), endowing its fine-grained position control capability. By decoupling instance-level spatial embeddings from semantic identity features, our method enables independent, accurate placement of each subject while introducing no extra train-time parameters or inference overhead. Restricting reference features to user-specified regions further enhances identity fidelity and spatial precision, unlocking new applications of controllable image customization (Figure 1).
@@
  • We design an automatic framework for data synthesis to obtain high-fidelity paired data, addressing the lack of precise subject consistency and positional control signals in existing public datasets.
  • We propose PositionIC, a lightweight position-controllable framework that decouples layout from subject. It achieves precise placement of multiple subjects through regional attention. At the same time, it explicitly enhances the subject fidelity of Diffusion transformers.
  • Extensive experimental results demonstrate that our method not only achieves the state-of-the-art performances on image customization, but also exhibits the highest precision control capability.
@@
Figure 3: The overall framework of PositionIC. (a) Reference images and prompts are encoded and concatenated with the latent embeddings $z_{t}$, then the whole token sequence is passed to the double stream blocks of DiT. (b) The mechanism of attention horizon. Each reference image $z_{r}^{i}$ is only visible for the specific area of latent noise $z_{t}$ in the attention map.
Related Work
Subject-driven Generation
In addition to using text prompts for conditional image generation, current diffusion models (Wei et al. 2023; Li, Li, and Hoi 2023; Huang et al. 2024; Ye et al. 2023; Xiong et al. 2025; Xiao et al. 2023; Bar-Tal et al. 2023; Feng et al. 2025) support reference image input to achieve preservation of subject identity. Dreambooth (Ruiz et al. 2023) and LoRa (Hu et al. 2022) control the generation of diffusion models through fine-tuning on the specific subject. Recently, Diffisuion-Transformers-based subject-driven models (Wu et al. 2025; Tan et al. 2024, 2025; Mou et al. 2025; Xiao et al. 2025; Labs et al. 2025) further advance subject-driven generation. They introduce reference images as context information through token concatenation to ensure the consistency of objects during generation process.
Position Controllable Generation
Some works (Wang et al. 2024b; Chen et al. 2024; Wang et al. 2024a; Jiménez 2023; Bar-Tal et al. 2023; Shi et al. 2025) have attempted to generate with precise layout control. Gligen (Li et al. 2023) encodes Fourier embedding as grounding tokens to inject position information. MS-diffusion (Wang et al. 2024c) utilizes a grounding resampler correlating visual information with specific entities and spatial constraints. However, they still suffer from inconsistencies in vision and position, which hinder further application.
Method
Bidirectional Multi-dimensional Perception Data Synthesis (BMPDS)
Position-controllable image-driven customization requires high-fidelity paired data, featuring prominent subjects and high resolution with layout control signals. However, existing open-source datasets such as Subject200K (Tan et al. 2024) generate paired data using diptych images, which leads to object inconsistency issues and are limited by low resolution and the lack of positional information. We introduce BMPDS, a Bidirectional Multi-dimensional Perception Data Synthesis framework to tackle these limitations. We adopt a hierarchical generation-and-selection strategy to progressively improve data quality, gradually introducing single-image and multi-image pairs with spatial control information during generation. The overall framework is depicted in Figure 2.
Customized Data Paired Synthesis
@@
Multi-dimensional Perception Data Filter
Previous studies (Cao et al. 2024; He et al. 2025) have shown that MLLMs have limited capability in recognizing fine-grained details in images. Instead of directly feeding image pairs into the MLLMs for filtering, we establish a reliable system to achieve more accurate and efficient data filtering after data generation. Specifically, we divide the filtering process into three levels based on granularity. Firstly, we segment the subjects from data pairs and utilize CLIP-I (Radford et al. 2021) and DINO (Caron et al. 2021) scores $s_{v}$ to filter out images with significantly lower consistency. Subsequently, we pass two subjects’ images through MLLMs (e.g., GPT-4o), which directly gives a similarity score $s_{vlm}$ based on shape, color, and details. Lastly, we employ Describe Anything Model (DAM) (Lian et al. 2025), a description expert to obtain detailed textual descriptions for each subject. Given these textual description pairs, we instruct GPT-4o as a judge to autonomously select comparative features(e.g., color, shape) and assign multi-dimensional similarity scores $s_{ds}$.
@@
We apply the filter on PIC-400K to rank the pairs and filter out inconsistent pairs. The filtered dataset PIC-98k consists of 44k single-subject pairs and 54k multi-subject pairs. More examples of our PIC-98K are depicted in Appendices.
Regional Attention Horizon
To manage subject consistency generation position simultaneously (e.g., generating the subject in a specific location in the image), we introduce a light-weight approach named attention accumulating, which can unlock DiT’s ability of spatial control without extra training and inference cost. As shown in Figure 3, the attention map can be divided into four areas: text-text self attention, image-image self attention, text-image and image-text attention. If extending to single or multi-subject custom generation via concatenating method like (Wu et al. 2025; Tan et al. 2024; Labs et al. 2025; Zhang et al. 2025a), the attention map will expand to nearly three times its original size, making it harder for the model to focus on the corresponding regions.
To achieve effective attention accumulation, we explicitly define the area that can be focused on for each reference image, which we name Attention Horizon. Previous works (Zhang et al. 2025a; Chen et al. 2024; Zhang et al. 2025b) explore the effort of restricting the attention horizon between special words and noise, we extend it to subject-driven generation task to unlock the positional control ability of diffusion transformers (DiT).
@@
Since the global text prompt contains information about the entire image, generating a reference object in a specific region requires focused and singular attention. They only need to pay attention to the corresponding region while ignoring other irrelevant tokens. Therefore, a limited attention horizon is required. We use a binary attention mask to shield the regions that each token should not directly focus on. As shown in Figure 3, each reference image $z_{ref}$ has a restricted attention horizon, which blocks the attention between itself and other reference images. Moreover, only the area within the bounding box in the noise is visible for the specific reference image. The mechanism of attention horizon mask $M$ can be formulated as:
@@
where $BOX_{i}$ is the bounding box of $i^{th}$ reference subject in the noise and $z_{t}^{n}$ is the $n^{th}$ patch of noise. Thus, the computation of attention is derived as:
@@
Experiments
Implementations
Training Detail
@@
Evaluation Metrics
For subject-generation tasks, we evaluate subject similarity using CLIP-I, DINO-I on Dreambench (Ruiz et al. 2023). For text fidelity, we calculate CLIP-T scores, which measure cosine similarity between the text embedding and image embedding from CLIP.
For the position-guided task, we evaluate different methods on our proposed PositionIC-Bench. We use Vision-R1 (Huang et al. 2025) to determine the bounding box of the subject and calculate mIoU and AP scores with the label.
@@
RPF (Chen et al. 2024) 0.341 0.015 / 0.063 / 0.007 0.369 0.070 / 0.002 / 0.011
MS-Diffusion (Wang et al. 2024c) 0.501 0.097 / 0.329 / 0.075 0.421 0.028 / 0.146 / 0.005
Instance-Diffusion (Wang et al. 2024b) 0.789 0.593 / 0.683 / 0.632 0.799 0.497 / 0.699 / 0.546
Gligen (Li et al. 2023) 0.808 0.632 / 0.865 / 0.811 0.825 0.628 / 0.858 / 0.811
@@
Qualitative Result
We visualize the comparison results with the current state-of-the-art methods in Figure 4. Overall, our method surpasses all current methods in terms of visual effects. It can be seen that our method can still effectively follow the prompt while maintaining subject consistency, demonstrating higher text fidelity. Other methods either fail to follow complex instructions or cannot maintain consistency. (e.g., UNO and DreamO cannot reproduce the dog face, while SSR-Encoder fails to add the Santa hat.) The results of the third row also reveal that PositionIC has a great capability on patterns and text. Furthermore, PositionIC consistently produces images with a higher degree of naturalness and visual plausibility.

### v4

@@ Abstract
Recent subject-driven image customization excels in fidelity, yet fine-grained instance-level spatial control remains an elusive challenge, hindering real-world applications. This limitation stems from two factors: a scarcity of scalable, position-annotated datasets, and the entanglement of identity and layout by global attention mechanisms. To this end, we introduce PositionIC, a unified framework for high-fidelity, spatially controllable multi-subject customization. First, we present BMPDS, the first automatic data-synthesis pipeline for position-annotated multi-subject datasets, effectively providing crucial spatial supervision. Second, we design a lightweight, layout-aware diffusion framework that integrates a novel visibility-aware attention mechanism. This mechanism explicitly models spatial relationships via an NeRF-inspired volumetric weight regulation to effectively decouple instance-level spatial embeddings from semantic identity features, enabling precise, occlusion-aware placement of multiple subjects. Extensive experiments demonstrate PositionIC achieves state-of-the-art performance on public benchmarks, setting new records for spatial precision and identity consistency. Our work represents a significant step towards truly controllable, high-fidelity image customization in multi-entity scenarios. Code and data will be publicly released.
1 Introduction
Diffusion-based models have recently revolutionized visual synthesis, especially in text-to-image (T2I) generation, where they produce photorealistic images closely aligned with short textual prompts [sdxl, stablediffusion3, ldm, playground, flux]. Beyond generic generation, subject-driven image customization aims to integrate one or more reference objects into user-specified scenes under the guidance of a text prompt. Traditionally, this task emphasizes two core goals: (i) preserving the appearance identity of the reference object(s), and (ii) aligning the synthesis with the user’s textual description, where remarkable improvements in visual fidelity have been achieved by recent methods [uno, dreamo, omini_control, wang2025unicombine].
@@ 1 Introduction
However, through exploring real-world applications like e-commerce product display, storybook illustration, and interior design, we identified an additional critical dimension: fine-grained, flexible spatial control, encompassing aspects such as subject placement, scale, occlusion, and relative layout. In other words, while prior works focus on “what” to generate, practical usage demands where and how each subject appears. To explicitly address this, we adopt a refined task definition: subject-driven customization with both identity fidelity and instance-level spatial controllability. This is a natural extension of the established task, not a fundamentally different objective, but one that better matches application requirements.
Despite substantial progress in identity-preserving personalization methods, current approaches still struggle with flexible placement and layout control. Specifically, two coupled bottlenecks remain: (1) the scarcity of large-scale datasets annotated with explicit positional labels of multiple subjects, hindering the learning of spatial reasoning; and (2) the prevalent use of global-level attention mechanisms in diffusion transformers, which inadvertently entangle semantic identity and spatial layout, thereby failing to provide fine-grained, visibility-aware control. While a few recent works [ms-diffusion, gligen] have attempted layout control, they typically exhibit a trade-off between spatial precision and subject identity consistency, thus limiting their practical utility.
To this end, we introduce two tightly coupled components: (i) a scalable Bidirectional Multi-dimensional Perception Data Synthesis (BMPDS) pipeline that generates high-quality, multi-subject, position-annotated data; and (ii) PositionIC, a lightweight layout-aware diffusion framework incorporating a visibility-aware attention mechanism that enables precise, controllable placement of multiple subjects while maintaining identity fidelity.
Specifically, to overcome the data bottleneck, we devise an automatic pipeline that expands single-subject collections into scalable, high-quality multi-subject datasets annotated with explicit position masks. A hierarchical training schedule establishes a bidirectional generation paradigm, progressively moving from single-to-multi subject synthesis and back, so that resolution constraints are relaxed while subject drift is suppressed. Because of the inherent hallucination of Multi-modal Large Language Models (MLLMs) [han2024instinctive, bpo], we avoid direct visual comparisons: expert vision models first translate visual content into textual descriptions, after which MLLMs perform multi-dimensional consistency checks. This two-stage filtering markedly improves data reliability.
Built upon the BMPDS corpus, PositionIC directly integrates our visibility-aware attention mechanism into the diffusion transformer architecture. This mechanism is achieved mainly via an operation inspired by NeRF [nerf] termed Volumetric Weight Regulation, which explicitly constructs attention masks based on the likelihood of visibility derived from volumetric rendering principles. Unlike prevalent global-level attention, our visibility-aware attention mechanism effectively decouples instance-level spatial embeddings from semantic identity features. This design enables independent and accurate placement of each subject by explicitly modeling occlusion and perspective relationships, and crucially introduces no extra train-time parameters or inference overhead. Restricting reference features to user-specified regions further enhances identity fidelity and spatial precision, unlocking new applications of controllable image customization (Figure 1).
@@
  • We present BMPDS, the first large-scale automatic data‐synthesis framework for high-fidelity, position-annotated multi‐subject pairs, filling a critical gap in spatial supervision.
  • We propose PositionIC, a lightweight yet highly effective layout-aware diffusion framework. It integrates a novel visibility-aware attention mechanism into the diffusion transformer architecture. This design explicitly decouples spatial layout from semantic identity, enabling unprecedentedly accurate multi-subject placement and occlusion-aware rendering without additional training parameters or inference overhead.
  • Extensive experimental results demonstrate that our approach delivers state-of-the-art performance on subject-driven customization benchmarks, achieving the highest spatial precision and identity consistency to date.
@@
Figure 3: The overall framework of PositionIC.(a) Reference images and prompts are encoded and concatenated with the latent embeddings $z_{t}$, then the whole token sequence is passed to DiT. Each reference image $z_{ref}^{i}$ is only visible for the specific area of latent noise $z_{t}$ in the attention map. (b) The objects’ binary mask and the semantic density $\sigma$ are used by VWR to calculate the weighted mask. The dashed boxes in the ray diagram represent the overlapping regions, where the same colors correspond to the weights in the mask.
2 Related Work
2.1 Subject-driven Generation
In addition to using text prompts for conditional image generation, current diffusion models [ELITE, li2023blip, huang2024realcustom, ip-adapter, groundingbooth, fastcomposer, MultiDiffusion, feng2025personalize] support reference image input to achieve preservation of subject identity. Dreambooth [dreambooth] and LoRa [lora] control the generation of diffusion models through fine-tuning on the specific subject. Recently, Diffisuion-Transformers-based subject-driven models [uno, omini_control, omini_control2, dreamo, xiao2025omnigen, flux-kontext] further advance subject-driven generation. They introduce reference images as context information through token concatenation to ensure the consistency of objects during generation.
2.2 Position Controllable Generation
Some works [instancediffusion, regional_flux, wang2024compositional, jimenez2023mixture, MultiDiffusion, shi2025layoutcot, LaRender] have attempted to generate with precise layout control. Gligen [gligen] encodes Fourier embedding as grounding tokens to inject position information. MS-diffusion [ms-diffusion] utilizes a grounding resampler correlating visual information with specific entities and spatial constraints. However, they still suffer from inconsistencies in vision and position, which hinder further application.
3 Method
3.1 Bidirectional Multi-dimensional Perception Data Synthesis (BMPDS)
Position-controllable image-driven customization requires high-fidelity paired data, featuring prominent subjects and high resolution with layout control signals. However, existing open-source datasets such as Subject200K [omini_control] generate paired data using diptych images, which leads to object inconsistency issues and are limited by low resolution and the lack of positional information. We introduce BMPDS, a Bidirectional Multi-dimensional Perception Data Synthesis framework to tackle these limitations. We adopt a hierarchical generation-and-selection strategy to progressively improve data quality, gradually introducing single-image and multi-image pairs with spatial control information during generation. The overall framework is depicted in Figure 2.
3.1.1 Customized Data Paired Synthesis
@@
3.1.2 Multi-dimensional Perception Data Filter
Previous studies [cao2024mmfuser, he2025analyzing] have shown that MLLMs have limited capability in recognizing fine-grained details in images. Instead of directly feeding image pairs into the MLLMs for filtering, we establish a reliable system to achieve more accurate and efficient data filtering. Specifically, we divide the filtering process into three levels based on granularity. Firstly, we segment the subjects from data pairs and utilize CLIP-I [CLIP] and DINO [DINO] scores $s_{v}$ to filter out images with significantly lower consistency. Subsequently, we pass two subjects’ images through MLLMs (e.g., GPT-4o), which directly gives a similarity score $s_{vlm}$ based on shape, color and so on. Lastly, we employ Describe Anything Model (DAM) [describe_anything], a description expert to obtain detailed textual descriptions for each subject. Given these textual description pairs, we instruct GPT-4o as a judge to autonomously select comparative features(e.g., color, shape) and assign multi-dimensional similarity scores $s_{ds}$.
@@ 3.1.2 Multi-dimensional Perception Data Filter
We apply the filter on PIC-400K to rank the pairs and filter out inconsistent pairs. The filtered dataset PIC-98k consists of 44k single-subject pairs and 54k multi-subject pairs. More examples of our PIC-98K are depicted in the appendix.
3.2 PositionIC
When extending text-to-image models to image-to-image tasks, the fidelity of the generated object largely depends on the magnitude of the attention between the corresponding regions in the target image and the reference image. We refer to this as attention accumulation. As shown in Figure 3, the attention map can be divided into four areas: text-text self attention, image-image self attention, text-image and image-text cross attention. If extending to single or multi-subject custom generation via concatenating method like [uno, omini_control, flux-kontext, eligen], the attention map will expand to nearly three times its original size, making it harder for the model to focus on the corresponding regions.
To achieve effective attention accumulation, we propose Visibility-Aware Attention, which utilizes volume rendering(VR) to explicitly define the area that can be focused on for each reference image. Previous works [eligen, regional_flux, zhang2025creatidesign, LaRender] explore the effort of restricting the attention horizon between special words and noise, we extend it to subject-driven generation task to unlock the positional control ability of diffusion transformers (DiT). The overall mechanism is shown in Figure 3.
3.2.1 Volume Rendering(VR)
The weighted mask utilizes VR [volume_render] to handle the occlusion relationships between individual objects. Similar to NeRF [nerf], the expected color $C(r)$ observed by the camera along the distance from the far bound $t_{f}$ to the near bound $t_{n}$ is:
$C(r)=\int_{t_{n}}^{t_{f}}T(t)\sigma(r(t))c(r(t),d)dt,$
$T(t)=exp(-\int_{t_{n}}^{t}\sigma(r(s))ds)$ (2)
where the camera ray $r(t)=o+td$, $\sigma(x)$ is volume density. We utilize the quadrature rule to obtain a discrete estimate of $C(\mathbf{r})$:
$\hat{C}(r)=\sum_{i=1}^{N}T_{i}(1-exp(-\sigma_{i}\delta_{i}))c_{i},$
$T_{i}=exp(-\sum_{j=1}^{i-1}\sigma_{i}\delta_{i})$ (3)
where $N$ denotes the spatial resolution for the distance range from near to far bounds. $\delta_{i}$ is the distance between adjacent samples.
3.2.2 Volumetric Weight Regulation
Similar to the 3D rendering process, controlling the generation position of individual objects on a 2D canvas can be viewed as a composite image captured by an orthographic virtual camera. To precisely distinguish the occlusion relationships between objects, we leverage VR to modulate a physically plausible semantic mask. Given the foreground mask $M_{i}$ and the sample distance $\delta$, the volumetric weight $\hat{M_{n}}$ can be formulate as:
$\hat{M}_{n}=exp(-\sum_{j=1}^{n-1}\sigma_{i}\delta)(1-exp(-\sigma_{n}\delta))M_{n}\hat{M}_{n-1}$ (4)
unlike in  Sec. 3.2.1, $\hat{M_{n}}$ is defined as the volumetric weight for each distance level, which is utilized to represent the degree of attention the image pays to each region of the reference image during attention calculation. Specifically, we utilize a semantic density $\sigma_{i}$ instead of the volume density to represent the semantic interaction between objects. Setting of $\sigma_{i}$ can be found in appendix.
3.2.3 Visibility-Aware Attention
@@
Since the global text prompt contains information about the entire image, generating a reference object in a specific region requires focused and singular attention. They only need to pay attention to the corresponding region while ignoring other irrelevant tokens. Therefore, we use a VWR-based attention mask, which we call visibility-aware attention(VAA), to shield the regions that each token should not directly focus on. As shown in Figure 3, each reference image $z_{ref}$ has a restricted attention horizon, which blocks the attention between itself and other reference images. Moreover, only the area within the bounding box in the noise is visible for the specific reference image. The mechanism of VAA mask $M$ can be formulated as:
@@
where $\hat{M}_{i}$ is volumetric weight of $i^{th}$ reference subject. Thus, the computation of attention is derived as:
@@ 3.3 PositionIC-Bench
4 Experiments
4.1 Implementations
4.1.1 Training Detail
@@
4.1.2 Evaluation Metrics
For subject-generation tasks, we evaluate subject similarity using CLIP-I, DINO-I on Dreambench [dreambooth]. For text fidelity, we calculate CLIP-T scores, which measure cosine similarity between the text embedding and image embedding from CLIP.
For the position-guided task, we evaluate different methods on our proposed PositionIC-Bench. We use Vision-R1 [visionr1] to determine the bounding box of the subject and calculate mIoU and AP scores with the label.
@@ 4.1.2 Evaluation Metrics
RPF [regional_flux] 0.341 0.015 / 0.063 / 0.007 0.369 0.070 / 0.002 / 0.011
MS-Diffusion [ms-diffusion] 0.501 0.097 / 0.329 / 0.075 0.421 0.028 / 0.146 / 0.005
Instance-Diffusion [instancediffusion] 0.789 0.593 / 0.683 / 0.632 0.799 0.497 / 0.699 / 0.546
Gligen [gligen] 0.808 0.632 / 0.865 / 0.811 0.825 0.628 / 0.858 / 0.811
@@
4.2 Qualitative Result
We visualize the comparison results with the current state-of-the-art methods in Figure 4. Overall, our method surpasses all current methods in terms of visual effects. It can be seen that our method can still effectively follow the prompt while maintaining subject consistency, demonstrating higher text fidelity. Other methods either fail to follow complex instructions or cannot maintain consistency. (e.g., UNO and DreamO cannot reproduce the dog face, while SSR-Encoder fails to add the Santa hat.) The results of the third row also reveal that PositionIC has a great capability on patterns and text. Furthermore, PositionIC consistently produces images with a higher degree of naturalness and visual plausibility. More examples can be found in Appendix.
@@ 4.2 Qualitative Result
4.3 Quantitative Evaluations
4.3.1 Subject-driven Analyses
Specifically, we discover that different object sizes can lead to inaccurate scores, which will be detailed in Appendix.